OpenStudy (anonymous):

Is the int of tsec^(2)2tdt = (t/2)tan(2t) - (1/4)log(sec2t) ?

OpenStudy (anonymous):

Series of explanation of the integral at x = 0 {"(2 t^3)/3+(2 t^3 x^2)/3+(4 t^3 x^4)/9+(34 t^3 x^6)/135+(124 t^3 \ x^8)/945+(2764 t^3 x^10)/42525+O(x^11)"}

OpenStudy (anonymous):

What? So \[\int\limits_{}^{}t \sec ^{2}2t\] = {"(2 t^3)/3+(2 t^3 x^2)/3+(4 t^3 x^4)/9+(34 t^3 x^6)/135+(124 t^3 \ x^8)/945+(2764 t^3 x^10)/42525+O(x^11)"} ???

OpenStudy (mathmate):

Hint: use integration by parts.

OpenStudy (anonymous):

I did...and the answer i got is \[(t/2) \tan2t - (1/4)logsec2t + C\]

OpenStudy (mathmate):

That's correct! If you are not sure, differentiate it with respect to t, you'll get back t*sec^2(2t). The other solution is a series expansion solution, which is also valid in cases where it is difficult to integrate analytically.

OpenStudy (anonymous):

u = t du= dt dv = sec^(2)t v = (1/2) tan2t

OpenStudy (anonymous):

I see...thank you!

OpenStudy (mathmate):

You're welcome! :)